40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 30     Entry time: Wed Oct 17 20:36:04 2012
Author: Koji 
Type: Optics 
Category: General 
Subject: RoC test cavity locked 

The RoC test setup has been built on the optical table at ATF.

The cavity formed by actual OMC mirrors have been locked.

The modulation frequency of the BB EOM was swept by the network analyzer.
A peak at ~30MHz was found in the transfer function when the input beam was misaligned and clipping was introduced at the transmission PD.
Without either the misalignment or the clipping, the peak disappears. Also the peak requires these imperfections to be directed in the same way
(like pitch and picth, or yaw and yaw). This strongly suggests that the peak is associated with the transverse mode.

The peak location was f_HOM = 29.79MHz. If we consider the length of the cavity is L=1.20m, the RoC is estimated as

RoC = L / (1 - Cos[f_HOM/(c/2/L) * PI]^2)

This formula gives us the RoC of 2.587 m.

I should have been able to find another peak at f_FSR-f_TMS. In deed, there was the structure found at 95MHz as expected.
However, the peak was really weak and the location was difficult to determine as it was coupled with the signal from residual RFAM.

The particle level in the clean booth was occasionally measured. Every measurement showed "zero".

To be improved:

  • The trans PD is 1801 which was found in ATF with the label of the 40m. It turned out that it is a Si PD.
    I need to find an InGaAs PD (1811, 1611, or my BBPD) or increase the modulation, or increase the detected light level.
    (==> The incident power on 1810 increased. Oct 17)
  • The BS at the transmission is actually Y1-45P with low incident angle. This can be replaced by 50% or 30% BS to increase the light on the fast PD.
    (==> 50% BS is placed. Oct 17)
  • I forgot to put a 50ohm terminator for the BB EOM.
    (==> 50Ohm installed. Oct 17)
  • A directional coupler could be used for the BBEOM signal to enhance the modulaiton by 3dB.
  • The mode matching is shitty. I can see quite strong TEM20 mode.
  • Use the longer cavity? L=1.8m is feasible on the table. This will move the peak at 27MHz and 56MHz (FSR=83MHz). Very promising.
    (==> L=1.8m, peak at 27MHz and 56MHz found. Oct.17)
Attachment 1: detailed_RoC_setup.pdf  147 kB  | Hide | Hide all
detailed_RoC_setup.pdf detailed_RoC_setup.pdf
Attachment 2: omc_cav_lock_SCRN_SHOT.png  374 kB  | Hide | Hide all
omc_cav_lock_SCRN_SHOT.png
Attachment 3: Cav_scan_response_20121016.pdf  32 kB  Uploaded Wed Oct 17 22:34:12 2012  | Hide | Hide all
Cav_scan_response_20121016.pdf
ELOG V3.1.3-